# Augmentation categories in higher dimensions

### v3

@@ Abstract.
For an exact symplectic manifold $M$ and a Legendrian submanifold $\Lambda$ of the contactification $M\times\mathbb{R}$, we construct the augmentation category (over a field of characteristic 2), a unital $A_{\infty}$-category whose objects are augmentations of the Chekanov-Eliashberg differential graded algebra. This extends the construction of the augmentation category by Ng-Rutherford-Shende-Sivek-Zaslow to contact manifolds of dimension greater than 3.
@@ 1. Introduction
In [BC], Bourgeois-Chantraine constructed, for every compact Legendrian submanifold in a $1$-jet space, a category $\mathcal{A}ug_{-}$, which is a non-unital $A_{\infty}$-category whose objects are augmentations of the Chekanov-Eliashberg DGA, using algebraic methods. In [NRSSZ], Ng-Rutherford-Shende-Sivek-Zaslow constructed, for every Legendrian link $\Lambda\in\mathbb{R}^{3}$, a category $\mathcal{A}ug_{+}$, which is a unital $A_{\infty}$-category whose objects are augmentations of the Chekanov-Eliashberg DGA of $\Lambda$, using parallel copies of $\Lambda$. We extend the construction of $\mathcal{A}ug_{+}$ (over a field of characteristic 2) to any compact Legendrian submanifold $\Lambda$ of the contactification $V=M\times\mathbb{R}$, where $M$ is an exact symplectic manifold with finite geometry at infinity (the important case to keep in mind is when $V=T^{*}X\times\mathbb{R}$ is a $1$-jet space). In the case of $M=T^{*}\mathbb{R}$, our construction agrees with the construction in [NRSSZ]. Roughly, the main theorem of [NRSSZ] says that in for a Legendrian link $\Lambda\subset\mathbb{R}^{3}$, $\mathcal{A}ug_{+}(\Lambda)$ is equivalent to a certain category of sheaves with microlocal rank 1. This paper is a step in generalising the equivalence to all $1$-jet spaces.
Very roughly, $\mathcal{A}ug_{+}(V,\Lambda)$ is defined as follows: objects are augmentations of the CE-DGA of $\Lambda$, morphisms are spanned by Reeb chords from $\Lambda$ to a positive push off of $\Lambda$. The $A_{\infty}$-operation $m_{k}$ is defined using the CE-DGA of $k+1$ parallel copies of $\Lambda$.
@@ Acknowledgments
I would like to thank Lenhard Ng for suggesting this problem, and for many helpful discussions. I would like to thank my advisor Robert Lipshitz for his guidance. I would like to thank Ville Nordström for providing a crucial insight, and many helpful discussions. I would like to thank Tobias Ekholm, John Etnyre, and Ian Zemke for helpful discussions.
@@ 2.1. Legendrian contact homology
In this section we review some background in Legendrian contact homology. We refer the reader to [EES2] for details, and [EENS] for the another exposition on the treatment for disconnected Legendrian submanifolds.
Let $(M,d\lambda)$ be a $2n$-dimensional exact symplectic manifold with finite geometry at infinity (see [EES2]*Definition 2.1 for the notion of finite geometry at infinity). Let $V=M\times\mathbb{R}$. Equip $V$ with the standard contact form $\alpha=dz-\lambda$, where $z$ is the $\mathbb{R}$ component ($V$ is sometimes called the contactification of $M$). Let $\Pi:V\to M$ denote the Lagrangian projection.
With the contact form $\alpha$, the Reeb vector field $R_{\alpha}$ satisfies $R_{\alpha}=\partial_{z}$. So, given a Legendrian submanifold $\Lambda\subset V$, the Reeb chords of $\Lambda$ correspond to self intersections of $\Pi(\Lambda)$. We can perturb $\Lambda$ so that there are finitely many Reeb chords, and they have distinct projections onto $M$ (if the Legendrian $\Lambda$ satisfies this condition, we call it chord generic). We denote the set of Reeb chords by $\mathcal{R}$.
Let $\Lambda\subset V$ be a chord generic compact Legendrian submanifold with vanishing Maslov class. Let $\mathcal{R}$ denote the set of Reeb chords on $\Lambda$. We fix a point $p_{j}\in\Lambda_{j}$ on each component $\Lambda_{j}$ of $\Lambda$, and for each Reeb chord endpoint in $\Lambda_{j}$ we choose an endpoint path connecting the endpoint to $p_{j}$. Furthermore, for j=1,,rj=1,...,r, we choose paths $\gamma_{1j}$ in $M$ connecting $\Pi(p_{1})$ to $\Pi(p_{j})$ and symplectic trivialisations of $\gamma_{1j}^{*}TM$ in which the tangent space $\Pi(T_{p_{1}}\Lambda)$ corresponds to the tangent space $\Pi(T_{p_{j}})$; for $j=1$, $\gamma_{11}$ is the trivial path. For any $1\leq i,j\leq r$, we can then define $\gamma_{ij}$ to be the path $\gamma_{1i}\cup\gamma_{1j}$ joining $\Pi(p_{i})$ to $\Pi(p_{j})$, and $\gamma_{ij}^{*}TM$ inherits a symplectic trivialisation from the trivialisations for $\gamma_{1i}$ and $\gamma_{1j}$.
@@
    where $\mu$ denotes the Maslov index.
@@
It is proved in [EES2] that for generic admissible $\Lambda$ and $J$, $\mathcal{M}_{\bar{A}}(q_{0};q_{1},...,q_{k})$ is a transversely cut out manifold of dimension $|q_{0}|-\sum_{i=1}^{k}|q_{i}|-1$ (for the precise notion of generic admissible, see the discussion surrounding [EES2]*Proposition 2.2). Furthermore, it admits a compactification as a manifold with boundary with corners in which the boundary consists of broken disks. Consequently, if the dimension equals 0 then it is compact.
@@ 2.2. Quantum flow trees
The moduli space of partial flow lines $\Delta$ with a fixed limit is the stable manifold of $\lim_{t\to\infty}\Delta(t)$, which means that the formal dimension of the moduli space at a partial flow line $\Delta$ is
$\dim(\Delta)=\dim(X)-\operatorname{ind}(\lim_{t\to\infty}\Delta(t)).$
The dimension of a complete flow line $\Delta$ with fixed limits is
$\dim(\Delta)=\operatorname{ind}(\lim_{t\to-\infty}\Delta(t))-\operatorname{ind}(\lim_{t\to\infty}\Delta(t))-1.$
@@
Let $V=J^{1}(X)=T^{*}X\times\mathbb{R}$ be a 1-jet space. A graphical Legendrian submanifold is a Legendrian submanifold of the form $\Gamma_{df,X}=\{(x,d_{x}f,f(x))\}$ for some smooth function $f:X\to\mathbb{R}$.
Let $(M,d\lambda)$ be a $2n$-dimensional exact symplectic manifold with finite geometry at infinity. Let $V=M\times\mathbb{R}$. Equip $V$ with the standard contact form $\alpha=dz-\lambda$, where $z$ is the $\mathbb{R}$ component. Let $\Lambda\subset V$ be a Legendrian submanifold. Let $U_{\Lambda}\subset V$ be a small contact neighborhood of $\Lambda$. By the Legendrian neighborhood theorem, we identify $U_{\Lambda}$ with a neighborhood of the 0-section of $J^{1}\Lambda$. Let $f_{1},f_{2}:\Lambda\to\mathbb{R}$ be small smooth functions such that $f_{1}>f_{2}$. Let $\Lambda^{\prime}=\Gamma_{df_{1},\Lambda}\sqcup\Gamma_{df_{2},\Lambda}\subset U_{\Lambda}\subset V$. Perturb the $f_{i}$ so that $\Lambda^{\prime}$ is quantum flow tree-generic, in the sense of [EL]*section 3.3.3 (which implies that $\Lambda^{\prime}$ is generic admissible in the sense of [EES2]). Note that $\Gamma_{df_{1},\Lambda}$ and $\Gamma_{df_{2},\Lambda}$ are disjoint. Choose a Riemannian metric $g$ on $\Lambda$.
@@
Let $\Delta$ be a (partial or complete) flow line of $\nabla(f_{1}-f_{2})$. The $1$-jet lift of $\Delta$ is $\pi^{-1}(\operatorname{Im}(\Delta))$, where $\pi:\Lambda^{\prime}\to\Lambda$ is the trivial 2-fold covering map obtained by restricting the base projection $\pi:U_{\Lambda}\to\Lambda$.
@@
A quantum flow tree $u_{\Delta}$ on $\Lambda^{\prime}$ consists of a holomorphic disk $u$ with boundary on $\Lambda$ with 1 positive puncture, and a finite set of partial flow lines $\Delta=\left\{\Delta_{j}\right\}_{j=1}^{m}$ of $\nabla(f_{1}-f_{2})$ each with a special puncture $x_{j},j=1,\ldots,m$, such that $x_{j}\in\partial u,j=1,\ldots,m$. Then the special punctures subdivide the boundary of $\partial u$ into arcs, and we require that there is a lift of these arcs to $\Lambda^{\prime}$ which together with the 1-jet lifts of the flow lines in $\Delta$ form closed curve when projected to $M$. A complete flow line of $\nabla(f_{1}-f_{2})$ on $\Lambda$ is also called a quantum flow tree.
The formal dimension of a quantum flow tree $u_{\Delta}$ (that is not a complete flow line) is
(2.8) $\dim(u_{\Delta})=\dim(u)+\sum_{\Delta_{j}\in\Delta}(\dim(\Delta_{j})-n+1).$
The formal dimension of a complete flow line $\Delta$ is still
$\dim(\Delta)=\operatorname{ind}(\lim_{t\to-\infty}\Delta(t))-\operatorname{ind}(\lim_{t\to\infty}\Delta(t))-1.$
@@
If the neighborhood $U_{\Lambda}$ is contained in a small neighborhood of $\Lambda$, then for any $\epsilon>0$, there exists an almost complex structure $J$, regular with respect to holomorphic disks with one positive puncture of dimension $\leq 1$, such that: there is a one-to-one correspondence between rigid holomorphic disks with one positive puncture and boundary on $\Lambda^{\prime}$, and rigid quantum flow trees on $\Lambda^{\prime}$ with one positive puncture. Furthermore, the $1$-jet lift of quantum flow trees in an $\epsilon$-neighborhood of the boundary lift of the corresponding holomorphic disk.
@@
Next we record some facts about rigid disks on $\Lambda^{\prime}$ (equivalently quantum flow trees) with a negative puncture at a minimum of $f_{1}-f_{2}$.
@@
Note that short Reeb chords of $\Lambda$ correspond to critical points of the function $f_{1}-f_{2}$.
@@
A rigid holomorphic disk with a negative puncture at a short Reeb chord corresponding to a minimum (index 0 critical point) of $f_{1}-f_{2}$ corresponds to either:
@@
    a rigid complete flow line of $f_{1}-f_{2}$, or
@@
    $u_{\Delta}$ where $u$ is a constant disk at some Reeb chord $b$ of $\Lambda$, and $\Delta$ consists of a single partial flow line, flowing from an endpoint of $b$ to a minimum of $f_{1}-f_{2}$. Here, a constant disk means a holomorphic strip (that is, a holomorphic disk with 1 positive puncture and 1 negative puncture) that equals $b\times\mathbb{R}$.
@@
Next we let $f_{0}>f_{1}>f_{2}:\Lambda\to\mathbb{R}$ be small smooth functions. Perturb the $f_{i}$ so that $\Lambda^{\prime}$ is quantum flow tree-generic, in the sense of [EL]*section 3.3.3. Assume that $f_{0}-f_{1}<\epsilon$ for some small $\epsilon$. Let $\Lambda^{\prime\prime}=\Gamma_{df_{0},\Lambda}\sqcup\Gamma_{df_{1},\Lambda}\sqcup\Gamma_{df_{2},\Lambda}$.
We note that rigid holomorphic disks on $\Lambda^{\prime\prime}$ correspond to quantum flow trees $u_{\Delta}$ with disk $u$ with boundary on $\Lambda^{\prime}$ and $\Delta$ a set of partial flow lines of $f_{0}-f_{1}$ with the appropriate lifts. To see this, let $f_{1.5}:\Lambda\to\mathbb{R}$ be a smooth function such that $f_{0}>f_{1}>f_{1.5}>f_{2}$, and that $f_{1.5}-f_{2}<\epsilon$. Let
$\Lambda^{\prime\prime\prime}=\Gamma_{df_{0},\Lambda}\sqcup\Gamma_{df_{1},\Lambda}\sqcup\Gamma_{df_{1.5},\Lambda}\sqcup\Gamma_{df_{2},\Lambda}.$
Perturb the $f_{i}$ so that $\Lambda^{\prime\prime\prime}$ is quantum flow tree-generic. Then, in the setting of Theorem 2.9, take $\Lambda=\Gamma_{df_{1},\Lambda}\sqcup\Gamma_{df_{2},\Lambda}$, and we see that $\Lambda^{\prime\prime}$ is in a small neighborhood of $\Gamma_{df_{1},\Lambda}\sqcup\Gamma_{df_{2},\Lambda}$, and the projection $\Lambda^{\prime\prime}\to\Gamma_{df_{1},\Lambda}\sqcup\Gamma_{df_{2},\Lambda}$ is a trivial 2-fold cover. Applying Theorem 2.9, and discarding the quantum flow trees with boundary lifts involving $\Gamma_{f_{1.5},\Lambda}$, we see the correspondence.
@@
Rigid holomorphic disks on $\Lambda^{\prime\prime}$ with a negative puncture at a short Reeb chord corresponding to a minimum of $f_{0}-f_{1}$ correspond to either:
@@
    a rigid complete flow line of $f_{0}-f_{1}$, or
@@
    a quantum flow tree $u_{\Delta}$ where $u$ is a constant disk at some Reeb chord $b$ of $\Lambda^{\prime}$, and $\Delta$ consists of a single partial flow line, flowing from an endpoint of $b$ to a minimum of $f_{0}-f_{1}$.

### v4

@@ Abstract.
For an exact symplectic manifold $M$ and a Legendrian submanifold $\Lambda$ of the contactization $M\times\mathbb{R}$, we construct the augmentation category (over a field of characteristic 2), a unital $A_{\infty}$-category whose objects are augmentations of the Chekanov-Eliashberg differential graded algebra. This extends the construction of the augmentation category by Ng-Rutherford-Shende-Sivek-Zaslow to contact manifolds of dimension greater than 3.
@@ 1. Introduction
In [BC], Bourgeois-Chantraine constructed, for every compact Legendrian submanifold in a $1$-jet space, a category $\mathcal{A}ug_{-}$, which is a non-unital $A_{\infty}$-category whose objects are augmentations of the Chekanov-Eliashberg DGA, using algebraic methods. In [NRSSZ], Ng-Rutherford-Shende-Sivek-Zaslow constructed, for every Legendrian link $\Lambda\in\mathbb{R}^{3}$, a category $\mathcal{A}ug_{+}$, which is a unital $A_{\infty}$-category whose objects are augmentations of the Chekanov-Eliashberg DGA of $\Lambda$, using parallel copies of $\Lambda$. We extend the construction of $\mathcal{A}ug_{+}$ (over a field of characteristic 2) to any compact Legendrian submanifold $\Lambda$ of the contactization $V=M\times\mathbb{R}$, where $M$ is an exact symplectic manifold with finite geometry at infinity (the important case to keep in mind is when $V=T^{*}X\times\mathbb{R}$ is a $1$-jet space). In the case of $M=T^{*}\mathbb{R}$, our construction agrees with the construction in [NRSSZ]. Roughly, the main theorem of [NRSSZ] says that in for a Legendrian link $\Lambda\subset\mathbb{R}^{3}$, $\mathcal{A}ug_{+}(\Lambda)$ is equivalent to a certain category of sheaves with microlocal rank 1. This paper is a step in generalising the equivalence to all $1$-jet spaces.
Very roughly, $\mathcal{A}ug_{+}(V,\Lambda)$ is defined as follows: objects are augmentations of the Chekanov-Eliashberg DGA of $\Lambda$, morphisms are cochain complexes, which as graded vector spaces, are each spanned by Reeb chords from $\Lambda$ to a positive push off of $\Lambda$. The differential $m_{1}$, and also higher $A_{\infty}$-operations $m_{k}$, are defined using the Chekanov-Eliashberg DGA of $k+1$ parallel copies of $\Lambda$ equipped with augmentations.
@@ Acknowledgments
I would like to thank Lenhard Ng for suggesting this problem, and for many helpful discussions. I would like to thank my advisor Robert Lipshitz for his guidance. I would like to thank Ville Nordström for providing a crucial insight, and many helpful discussions. I would like to thank Tobias Ekholm, John Etnyre, and Ian Zemke for helpful discussions. I would like to thank the referee for the careful read of this paper and for helpful comments.
@@ 2.1. Legendrian contact homology
In this section we review some background in Legendrian contact homology. We refer the reader to [EES2] for details, and [EENS] for another exposition on the treatment for disconnected Legendrian submanifolds.
Let $(M,d\lambda)$ be a $2n$-dimensional exact symplectic manifold with finite geometry at infinity (see [EES2]*Definition 2.1 for the notion of finite geometry at infinity). Let $V=M\times\mathbb{R}$. Equip $V$ with the standard contact form $\alpha=dz-\lambda$, where $z$ is the $\mathbb{R}$ component ($V$ is sometimes called the contactization of $M$). Let $\Pi:V\to M$ denote the Lagrangian projection.
With the contact form $\alpha$, the Reeb vector field $R_{\alpha}$ satisfies $R_{\alpha}=\partial_{z}$. So, given a Legendrian submanifold $\Lambda\subset V$, the Reeb chords of $\Lambda$ correspond to self intersections of $\Pi(\Lambda)$. We can perturb $\Lambda$ so that there are finitely many Reeb chords, and the corresponding self intersections of $\Pi(\Lambda)$ are transverse double points (if the Legendrian $\Lambda$ satisfies this condition, we call it chord generic). We denote the set of Reeb chords by $\mathcal{R}$.
Let $\Lambda\subset V$ be a chord generic compact Legendrian submanifold with vanishing Maslov class. Let $\mathcal{R}$ denote the set of Reeb chords on $\Lambda$. We fix a point $p_{j}\in\Lambda_{j}$ on each component $\Lambda_{j}$ of $\Lambda$. We pick one of the basepoints, say $p_{1}$ for convenience, which we call the special basepoint. For each Reeb chord endpoint in $\Lambda_{j}$ we choose an endpoint path connecting the endpoint to $p_{j}$. Furthermore, for j=1,,rj=1,...,r, we choose paths $\gamma_{1j}$ in $M$ connecting $\Pi(p_{1})$ to $\Pi(p_{j})$ and symplectic trivialisations of $\gamma_{1j}^{*}TM$ in which the tangent space $\Pi(T_{p_{1}}\Lambda)$ corresponds to the tangent space $\Pi(T_{p_{j}})$; for $j=1$, $\gamma_{11}$ is the trivial path. For any $1\leq i,j\leq r$, we can then define $\gamma_{ij}$ to be the path $\gamma_{1i}\cup\gamma_{1j}$ joining $\Pi(p_{i})$ to $\Pi(p_{j})$, and $\gamma_{ij}^{*}TM$ inherits a symplectic trivialisation from the trivialisations for $\gamma_{1i}$ and $\gamma_{1j}$.
@@
    where $\mu$ denotes the Maslov index. Note that the grading is independent of choices of endpoint paths in the case where $\Lambda$ is connected, and the grading depends on the choices of paths in the case where $\Lambda$ is disconnected.
    In the case where $\Lambda$ is disconnected, and $M\times\mathbb{R}=J^{1}(X)$ is a $1$-jet space, the choice of these paths is equivalent to choosing a Maslov potential, but we will not need that in this paper. A Maslov potential on the $\Lambda$ assigns an integer to points on $\Lambda$, in a way that is locally constant, except for at cusps of the front projection, where the integer increases by one when going up through a cusp.
@@
It is proved in [EES2] that for generic admissible $\Lambda$ and $J$, $\mathcal{M}_{\bar{A}}(q_{0};q_{1},...,q_{k})$ is a transversely cut out manifold of dimension $|q_{0}|-\sum_{i=1}^{k}|q_{i}|-1$. Roughly, admissibility here means that around the endpoints of Reeb chords, $\Lambda$ is real analytic with respect to some standard coordinate given by $J$, and generic admissible means that we further perturb $\Lambda$ in the space of admissible Legendrian submanifolds (for the precise notion of generic admissible, see the discussion surrounding [EES2]*Section 2.3). Furthermore, it admits a compactification as a manifold with boundary with corners in which the boundary consists of broken disks. Consequently, if the dimension equals 0 then it is compact.
@@ 2.2. Quantum flow trees
Since a partial flow line is determined by its special puncture, we have that the moduli space of partial flow lines $\Delta$ with $\lim_{t\to\infty}\Delta(t)$ fixed is the stable manifold of $\lim_{t\to\infty}\Delta(t)$, which means that the formal dimension of the moduli space at a partial flow line $\Delta$ is
$\dim(\Delta)=\dim(X)-\operatorname{ind}\left(\lim_{t\to\infty}\Delta(t)\right).$
The dimension of the moduli space of complete flow lines $\Delta$, with $\lim_{t\to\pm\infty}\Delta(t)$ fixed, modulo the translational $\mathbb{R}$-action, is
$\dim(\Delta)=\operatorname{ind}\left(\lim_{t\to-\infty}\Delta(t)\right)-\operatorname{ind}\left(\lim_{t\to\infty}\Delta(t)\right)-1.$
@@
Let $V=J^{1}(X)=T^{*}X\times\mathbb{R}$ be a 1-jet space. A graphical Legendrian submanifold is a Legendrian submanifold of the form $\Gamma_{df,X}=\{\left(x,d_{x}f,f(x)\right)\}$ for some smooth function $f:X\to\mathbb{R}$.
Let $(M,d\lambda)$ be a $2n$-dimensional exact symplectic manifold with finite geometry at infinity. Let $V=M\times\mathbb{R}$. Equip $V$ with the standard contact form $\alpha=dz-\lambda$, where $z$ is the $\mathbb{R}$ component. Let $\Lambda\subset V$ be a Legendrian submanifold. Let $U_{\Lambda}\subset V$ be a small contact neighborhood of $\Lambda$. By the Legendrian neighborhood theorem, we identify $U_{\Lambda}$ with a neighborhood of the 0-section of $J^{1}\Lambda$. Let $f_{1},f_{2}:\Lambda\to\mathbb{R}$ be small smooth functions such that $f_{1}<f_{2}$. Let $\Lambda^{\prime}=\Gamma_{df_{1},\Lambda}\cup\Gamma_{df_{2},\Lambda}\subset U_{\Lambda}\subset V$. Note that $\Gamma_{df_{1},\Lambda}$ and $\Gamma_{df_{2},\Lambda}$ are disjoint. Choose a Riemannian metric $g$ on $\Lambda$. Perturb the $f_{i},g$ so that $\Lambda^{\prime}$ is quantum flow tree-generic, in the sense of [EL]*Section 3.3.3, which implies that $\Lambda^{\prime}$ is generic admissible in the sense of [EES2]*Sections 2.3.
Roughly, the data of an almost complex structure, with $\Lambda$, with $f_{2}-f_{1}$ is quantum flow tree-generic in the sense of [EL]*Section 3.3.3, if:
  • the moduli spaces of Morse flow lines of $f_{2}-f_{1}$ are transversely cut out,
  • the almost complex structure with $\Lambda$ is generic admissible,
  • partial flow trees are transverse to the boundary evaluation maps of the transversely cut out holomorphic curves.
@@
Let $\Delta$ be a (partial or complete) flow line of $\nabla(f_{2}-f_{1})$. The $1$-jet lift of $\Delta$ is $\pi^{-1}(\operatorname{Im}(\Delta))$, where $\pi:\Lambda^{\prime}\to\Lambda$ is the trivial 2-fold covering map obtained by restricting the base projection $\pi:U_{\Lambda}\to\Lambda$.
@@
A quantum flow tree $u_{\Delta}$ on $\Lambda^{\prime}$ consists of a punctured holomorphic disk $u:(D_{k}^{2},\partial D_{k}^{2})\to(V\times\mathbb{R},\Lambda\times\mathbb{R})$ with 1 positive puncture satisfying the same asymptotic conditions as the holomorphic disks in the differential of the Chekanov-Eliashberg DGAs, and a finite set of partial flow lines $\Delta=\left\{\Delta_{j}\right\}_{j=1}^{m}$ of $\nabla(f_{2}-f_{1})$ each with a special puncture $x_{j}$, such that $x_{j}\in\partial\tilde{u}$ for all $j$. Here, $\tilde{u}:(D_{k}^{2},\partial D_{k}^{2})\to(V\times\mathbb{R},\Lambda\times\mathbb{R})\to(V,\Lambda)$ is given by $u$ composed with the projection map. Then the special punctures subdivide the boundary $\partial\tilde{u}$ into arcs on $\Lambda$. For these arcs on $\Lambda$, we pick a $1$-jet lift to $\Lambda^{\prime}$. More specifically, a $1$-jet lift consists of arcs on $\Lambda^{\prime}$, and the projection map $\pi:\Lambda^{\prime}\to\Lambda$ restricted to the $1$-jet lift is a homeomorphism onto the original arcs on $\Lambda$. The 1-jet lift of these arcs, together with the 1-jet lifts of the flow lines in $\Delta$, are required to form a closed curve when projected to $M$. We say that the 1-jet lift of a quantum flow tree is the union of these arcs and the 1-jet lifts of the flow lines in $\Delta$. The 1-jet lift of a quantum flow tree is considered part of the information of a quantum flow tree. A complete flow line of $\nabla(f_{2}-f_{1})$ on $\Lambda$ is also considered a quantum flow tree.
By intersecting the moduli spaces of special punctures (which are isomorphic to the moduli spaces of the partial flow lines) with the boundary arcs of the moduli space of $u$, we see that the formal dimension of a quantum flow tree $u_{\Delta}$ (that is not a complete flow line), up to domain reparametrisation and $\mathbb{R}$-translation of $u$, is
(2.8) $\dim(u_{\Delta})=\dim(u)+\sum_{\Delta_{j}\in\Delta}(\dim(\Delta_{j})-n+1),$
where $\dim(u)$ is the dimension of the moduli spaces of holomorphic disks with the same boundary conditions as $u$, up to domain reparametrisation and $\mathbb{R}$-translation. A version of the above formula for the $n=2$ case appears in [EENS]*Section 4.1, for multiscale flow trees, which are similar and related to quantum flow trees. The formal dimension of a complete flow line $\Delta$ up to the $\mathbb{R}$-translation is still
$\dim(\Delta)=\operatorname{ind}\left(\lim_{t\to-\infty}\Delta(t)\right)-\operatorname{ind}\left(\lim_{t\to\infty}\Delta(t)\right)-1.$
@@
If the neighborhood $U_{\Lambda}$ is contained in a small neighborhood of $\Lambda$, then for any $\epsilon>0$, there exists an almost complex structure $J$ on $V\times\mathbb{R}$, compatible with the symplectisation, and regular with respect to holomorphic disks with one positive puncture of dimension $\leq 1$, such that there is a one-to-one correspondence between:
  • rigid holomorphic disks $(D_{k}^{2},\partial D_{k}^{2})\to(V\times\mathbb{R},\Lambda^{\prime}\times\mathbb{R})$ with one positive puncture satisfying the same asymptotic conditions as the holomorphic disks in the differential of the Chekanov-Eliashberg DGAs,
  • and rigid quantum flow trees on $\Lambda^{\prime}$.
Furthermore, the $1$-jet lift of a quantum flow tree lies in an $\epsilon$-neighborhood of the boundary of the corresponding holomorphic disk (after we project the boundary of the corresponding holomorphic disk to $\Lambda$).
@@
Next we record some facts about rigid disks on $\Lambda^{\prime}$ (equivalently quantum flow trees) with a negative puncture at a minimum of $f_{2}-f_{1}$.
@@
Note that short Reeb chords of $\Lambda$ correspond to critical points of the function $f_{2}-f_{1}$.
@@
A rigid holomorphic disk with a negative puncture at a short Reeb chord corresponding to a minimum (index 0 critical point) of $f_{2}-f_{1}$ corresponds to either:
@@
    a rigid complete flow line of $f_{2}-f_{1}$, or
@@
    $u_{\Delta}$ where $u$ is a constant disk at some Reeb chord $b$ of $\Lambda$, and $\Delta$ consists of a single partial flow line, flowing from an endpoint of $b$ to a minimum of $f_{2}-f_{1}$. Here, a constant disk means a holomorphic strip (that is, a holomorphic disk with 1 positive puncture and 1 negative puncture) that equals $b\times\mathbb{R}$.
@@
Next we let $f_{1}<f_{2}<f_{3}:\Lambda\to\mathbb{R}$ be small smooth functions. Perturb the $f_{i}$ so that $\Lambda^{\prime}$ is quantum flow tree-generic, in the sense of [EL]*section 3.3.3. Assume that $\left\|f_{2}-f_{1}\right\|_{C^{2}}<\epsilon$ for some small $\epsilon$. Let $\Lambda^{\prime\prime}=\Gamma_{df_{1},\Lambda}\cup\Gamma_{df_{2},\Lambda}\cup\Gamma_{df_{3},\Lambda}$.
We note that rigid holomorphic disks on $\Lambda^{\prime\prime}$ correspond to quantum flow trees $u_{\Delta}$ with disk $u$ with boundary on $\Lambda^{\prime}$ and $\Delta$ a set of partial flow lines of $f_{2}-f_{1}$ with the appropriate lifts. To see this, let $f_{2.5}:\Lambda\to\mathbb{R}$ be a smooth function such that $f_{1}<f_{2}<f_{2.5}<f_{3}$, and that $\left\|f_{3}-f_{2.5}\right\|_{C^{2}}<\epsilon$. Let
$\Lambda^{\prime\prime\prime}=\Gamma_{df_{1},\Lambda}\cup\Gamma_{df_{2},\Lambda}\cup\Gamma_{df_{2.5},\Lambda}\cup\Gamma_{df_{3},\Lambda}.$
Perturb the $f_{i}$ so that $\Lambda^{\prime\prime\prime}$ is quantum flow tree-generic. Then, in the setting of Theorem 2.9, take $\Lambda=\Gamma_{df_{2},\Lambda}\cup\Gamma_{df_{3},\Lambda}$, and we see that $\Lambda^{\prime\prime\prime}$ is in a small neighborhood of $\Gamma_{df_{2},\Lambda}\cup\Gamma_{df_{3},\Lambda}$, and the projection $\Lambda^{\prime\prime\prime}\to\Gamma_{df_{2},\Lambda}\cup\Gamma_{df_{3},\Lambda}$ is a trivial 2-fold cover. Applying Theorem 2.9, and discarding the quantum flow trees with boundary lifts involving $\Gamma_{f_{2.5},\Lambda}$, we see the correspondence. By Lemma 2.11, we also have the following:
@@
Rigid holomorphic disks on $\Lambda^{\prime\prime}$ with a negative puncture at a short Reeb chord corresponding to a minimum of $f_{2}-f_{1}$ correspond to either:
@@
    a rigid complete flow line of $f_{2}-f_{1}$, or
@@
    a quantum flow tree $u_{\Delta}$ where $u$ is a constant disk at some Reeb chord $b$ of $\Lambda^{\prime}$, and $\Delta$ consists of a single partial flow line, flowing from an endpoint of $b$ to a minimum of $f_{2}-f_{1}$.
Since Corollary 2.12 allows us to compute some holomorphic disks on $\Lambda^{\prime\prime}=\Gamma_{df_{1},\Lambda}\cup\Gamma_{df_{2},\Lambda}\cup\Gamma_{df_{3},\Lambda}$ by considering holomorphic disks on $\Gamma_{df_{2},\Lambda}\cup\Gamma_{df_{3},\Lambda}$ and flow lines of $f_{2}-f_{1}$, we say that we degenerate $\Gamma_{df_{1},\Lambda}\cup\Gamma_{df_{2},\Lambda}\cup\Gamma_{df_{3},\Lambda}$ onto $\Gamma_{df_{2},\Lambda}\cup\Gamma_{df_{3},\Lambda}$.